OpenStudy (anonymous):

what does the little degree sign mean, in functions? I see it in my algebra and I know it's a composite function but I can never remember what it is I'm supposed to do...

Nnesha (nnesha):

are you talking about ( f o g)(x) that just means function in a function...

Nnesha (nnesha):

for example if it is f o g(2) so just asking to find f(g(2)) ..

OpenStudy (anonymous):

ohhh so it's multiplying the two functions? or is it plugging in 2 (in that case) for x?

Nnesha (nnesha):

\[\large\rm g(x) =x+2 ~~~~~~~~~~~~f(x)=4x+2\] find( f o g)(2) first substitute 2 into g(X) function and then substitute the result into f(x) \[g(2)=2+2=\color{Red}{4}\] \[\large\rm (f o g)(2) \rightarrow f(\color{Red}{g(2)}) = 4(\color{Red}{4})+2\]

OpenStudy (jdoe0001):

\(\bf f(x)=2x^2+5-3x\qquad {\color{brown}{ g(x)}} =meow \\ \quad \\ (f\circ g)(x)\implies f(\quad {\color{brown}{ g(x)}}\quad )=2{\color{brown}{ (meow)}}^2+5-3{\color{brown}{ (meow)}}\)

OpenStudy (anonymous):

ohhhhhhhhhhhhh that's makes sense now omg lol

OpenStudy (anonymous):

thanks guys!

Nnesha (nnesha):

substitution not multiplication

OpenStudy (anonymous):

right, plugging numbers in

Nnesha (nnesha):

yep.

Nnesha (nnesha):

yw!